This Question: 1 pt This Quiz: 10 pts Question Help na□ (Round up to the nearest observation.) empts 1 of 2 0 of 0 of Enter your answer in the answer box. 0 of o 2 6
Answer #1

Solution:- Given that limits (0.45,0.75)

p = (upper + lower)/2 = (0.75+0.45)/2 = 0.6

E = (upper - lower)/2 = (0.75-0.45)/2 = 0.15


95% confidence interval for Z = 1.96

sample n = (Z/E)^2*p*q = (1.96/0.15)^2*0.6*0.4 = 40.977 = 41


n = 41